• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:03:17 on February25,2002

    Entry number 4195

    keyword=Shift Summary

    Day shift:
    John LeRose (SL)
    G. Kumbartzki (TO)
    G. Laveissiere (3rd)
    Program is continue to take data.
    Plagued the entire shift by annoying alarms for too high positive low voltage on the left VDC. (AAAAARGHHHHHHH!)
    run ....start
    3137 07:57
    3138 08:10
    3139 08:22
    3140 08:39
    3141 08:51

    Cryogenics problem believed to have originated at the CHL causes magnets to trip at 9:05
    Lost Left Dipole, left Q1, and the whole right arm
    Access required to reset power supplies.
    09:40 all power supplies reset and magnets ramping back up.
    09:55 Beam back
    While waiting for magnets to settle did target thickness vs beam current studies

    3142 10:07 FPP trip at 10:13 Four minutes to reset
    3143 10:20 FPP trip at 10:22 Four minutes to reset
    3144 10:32
    3145 10:45
    3146 10:57 FPP trip at 11:08
    Beginning of Computer deadtime problems
    3147 11:09
    3148 11:27
    3149 11:48
    3150 12:02
    Had problems with excessive and erratic computer deadtime (20-80%). Killed CODA and reset everything. Didn't help. Called Bob
    3151 12:12
    3152 12:20
    3153 12:27
    3154 12:30
    3155 12:38 at 12:52 FPP trip restored by 12:58
    3156
    3157 12:57
    Computer deadtime problems persist
    13:12 called for controlled access to fix computer deadtime (bad fastbus module?)
    14:08 Bob says the problem is solved. Coming out of access. See Halog 4171
    14:31 beam is back
    runs 3157-3163 Diagnostics/junk
    3164 14:42
    3165 14:54
    Target IOC rebooted at 15:00
    3166 15:10
    3167 15:22
    3168 15:35
    3169 15:47
    3170 having roc14 problems at end of shift